Bayside insurance offers two health plans. Under plan a, Giselle would have to pay the first $70 of her medical bills, plus 35% of the rest. Under plan b, Giselle would have to pay the first $230 and 30% for of the rest. For what amount of medical bills will plan b save Giselle money?
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

let the bill amount be x
Plan A= 75 + 35.00% of rest
Plan B= 230 + 30.00% of the rest
Plan B < Plan A
230+0.35(x-230)<75+0.35*x
230+0.30x-80.5<75+0.35x-26.25
230-75-80.5+26.25<0.35x-0.3x
100.75<0.05x
2015 the bill should be greater than $2015

Answer by ikleyn(53742)

.
Bayside insurance offers two health plans. Under plan A, Giselle would have to pay the first $70 of her medical bills,
plus 35% of the rest.
Under plan B, Giselle would have to pay the first $230 and 30% for of the rest.
For what amount of medical bills will plan B save Giselle money?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        In the post by @mananth,  the conception,  the numbers,  the solution and the answer are incorrect.
        I came to bring a correct solution.


let the bill amount be 	x							

    Plan A =  70 + 0.35*(x - 70).

    Plan B = 230 + 0.3*(x - 230).					


Plan B < Plan A	
						
    230 + 0.3*(x-230) < 70 + 0.35*(x-70)

    230 + 0.30x - 69 < 70 + 0.35x - 24.5

    230 - 69 - 70 + 24.5 < 0.35x - 0.3x	

    115.5 < 0.05x 

    X > 115.5/0.05 = 2310


ANSWER.  Plan B is cheaper if the bill is greater than $2310.

Solved correctly.